OpenStudy (anonymous):

Find the vertex f(x)= 3x^2-18x+3 please help me. Awards will be given

OpenStudy (anonymous):

use the all mighty formula \[x= -\frac{b}{2a}\] to find the first coordinate

OpenStudy (anonymous):

in this case \[a=3,b=-18\] so \[-\frac{b}{2a}=-\frac{-18}{2\times3}=3\]

OpenStudy (anonymous):

that is the first coordinate and the second coordinate is what you get when you replace x by 3. you get \[3\times 3^2-18\times 3+3=27-54+3=-27+3=-24\] check my arithmetic because it is late

OpenStudy (anonymous):

also because i didn't really compute. as soon as i saw 27 i knew the next one would be -54. always works that way

OpenStudy (anonymous):

vertex is (3,-24) unless i messed up

OpenStudy (anonymous):

so (3,-24) ? and that would make it a maximum right?

OpenStudy (anonymous):

heck no

OpenStudy (anonymous):

this is a parabola that faces up. no max, but a min

OpenStudy (anonymous):



OpenStudy (anonymous):

the minimum value is -24

OpenStudy (anonymous):

Ah I got it mixed up!

OpenStudy (anonymous):

I always do that with the minimum maximum.